Convert the vector A = xy ax - 2x ay to cylindrical system.



1
Expert's answer
2021-02-08T18:40:08-0500

Given vector,

"A=xy(a_x-2xa_y)"


"\Rightarrow A= xya_x-2x^2ya_y"

Now, "x=a\cos\theta"

"y=a\sin\theta"

"z=z"

"B=a\cos\theta\times a\sin\theta\times a_y-2\times (a\cos\theta)^2\times a\sin\theta\times a_y"

we know that "a=a_x= a_y" for the circle

"=a^3(cos\theta.\sin\theta-2\cos^2\theta.\sin\theta )"


"=a^3(\frac{1}{2}\sin2\theta-cos\theta. \sin2\theta)"


"=a^3\sin2\theta(\frac{1-2\cos\theta}{2})"


"=\frac{4a^3\sin2\theta \cos^2(\frac{\theta}{2})}{2}"


"=2a^3\sin2\theta \cos^2(\frac{\theta}{2})"